Question: The sum of two digits of a number is 10. If the digits chane their places, the value of the number is decreased by 18. Find the number.
Answer by Mathtut(3670)

let the number be ab.....which can be written as 10a+b......the number reversed would be ba.....which can be written as 10b+a
:
a+b=10.......eq 1
10b+a=10a+b-18...eq 2--->9b-9a=-18
:
rewrite eq 1 to a=10-b and plug it into eq 2
:
9b-9(10-b)=-18
:
9b-90+9b=-18
:
18b=72
:
b=4
:
a=10-4 =6
:
so our number is 64
: